OpenStudy (anonymous):

Find the Vertex and all Intercepts of the parabola. f(x)=2x^2-4x This is a practice one; however I am having trouble understanding how to get to the point where I need to sketch it on the graph and determine my values..

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

And are you in a calculus course?

OpenStudy (anonymous):

Not really...I've been looking at this and i'm completely lost. Strange, but when i get to this type of problem, I go blank. Obviously I can graph it, but can't figure out how to solve it.

OpenStudy (anonymous):

Well finding the intercepts is easy enough. For the x-intercepts we will set y= 0 For the y-intercepts we will set x=0 For the vertex read this: http://www.purplemath.com/modules/sqrvertx.htm

OpenStudy (anonymous):

y = x(2x - 4) set y = 0 which means either x = 0 or 2x-4=0 x=2,

OpenStudy (anonymous):

Tell me if you need clarification on anything! =D

OpenStudy (anonymous):

Ill look at that link to see if I can figure out the vertex Thank you for your help